# Neutrino Constraints on Black Hole Formation in M31

### v1

@@ 5 Summary
We presented a detailed comparison of neutrino emission predictions across various equations of state (EOS), including Lattimer-Swesty (LS180, LS220), Shen (Shen-TM1e, Shen-TM1), Togashi, and SFHo. Results clearly illustrate the strong dependence of neutrino emission on EOS choice, with models employing stiff EOS and higher progenitor masses generally predicting more energetic neutrino outputs.
Using these theoretical predictions, we estimated the expected neutrino detection numbers at Super-Kamiokande (SK) for a collapse event occurring at the distance of M31 ($\sim$770 kpc). The observed non-detection of neutrino events around the time of M31-2014-DS1 places meaningful constraints on model parameters. Specifically, parameter combinations predicting higher neutrino energies and fluxes can be excluded at high confidence levels, thus providing valuable feedback to theoretical modeling efforts.
@@ Acknowledgments
This work is supported by JSPS KAKENHI grant Nos. JP21K13913, JP23KJ2150, JP24H02236, JP24H02245, JP24K00632, JP24K00668, JP24K07021 JP25K01035, and JP25K17399. Numerical computations were, in part, carried out on a computer cluster at CfCA of the National Astronomical Observatory of Japan. This work was supported in part by MEXT as “Program for Promoting Research on the Supercomputer Fugaku” (Toward a unified view of the universe: from large scale structures to planets) and, in part by the Inter-University Research Program of the Institute for Cosmic Ray Research (ICRR), the University of Tokyo.

### v2

@@ 5 Summary
We presented a detailed comparison of neutrino emission predictions across various equations of state (EOS), including Lattimer-Swesty (LS180, LS220), Shen (Shen-TM1e, Shen-TM1), Togashi, DD2F RDF, and SFHo. Results clearly illustrate the strong dependence of neutrino emission on EOS choice, with models employing stiff EOS and higher progenitor masses generally predicting more energetic neutrino outputs.
Using these theoretical predictions, we estimated the expected neutrino detection numbers at Super-Kamiokande (SK) for a collapse event occurring at the distance of M31 ($\sim$770 kpc). The observed non-detection of neutrino events around the time of M31-2014-DS1 places meaningful constraints on model parameters. Specifically, parameter combinations predicting higher neutrino energies and fluxes can be excluded at high confidence levels, thus providing valuable feedback to theoretical modeling efforts. In this analysis, we did not consider the details of event selection such as energy cuts, in order to keep the estimate simple. When applying to real data, various cut conditions need to be taken into account, which would generally relax the constraints compared to those presented here (Nakanishi et al., 2025).
@@ Acknowledgments
We thank NK Largani and T. Fischer for providing their data. This work is supported by JSPS KAKENHI grant Nos. JP21K13913, JP23KJ2150, JP24H02236, JP24H02245, JP24K00632, JP24K00668, JP24K07021 JP25K01035, and JP25K17399. Numerical computations were, in part, carried out on a computer cluster at CfCA of the National Astronomical Observatory of Japan. This work was supported in part by MEXT as “Program for Promoting Research on the Supercomputer Fugaku” (Toward a unified view of the universe: from large scale structures to planets) and, in part by the Inter-University Research Program of the Institute for Cosmic Ray Research (ICRR), the University of Tokyo.